Solve the quadratic equation by factorization.
X²-64=0
could I be shown it by step by step please

Answer by Cromlix(4381)

Hi there,
x^2 - 64 = 0
This is referred to as a difference of squares.
It factorises to:-
(x - 8)(x + 8)= 0
x - 8 = 0
x = 8
x + 8 = 0
x = -8
The two values for x are
8 and -8
.........
To spot another difference of squares
look for x^2 at the beginning. Followed by
a negative sign. Then an integer able to
be square rooted to a whole number.
Examples:
x^2 - 169 = 0
(x - 13)(x + 13)= 0
4x^2 - 36 = 0
(2x - 6)(2x + 6) = 0
Hope this helps:-)